# Tau Tridents at Accelerator Neutrino Facilities

### v1

@@ Abstract
We present the first detailed study of Standard Model (SM) neutrino tridents involving tau leptons at the near detectors of accelerator neutrino facilities. These processes were previously thought to be negligible, even at future facilities like DUNE, based on approximations that underestimated the tau trident cross sections. Our full calculation, including both coherent and incoherent scatterings, reveals that the DUNE near detector will actually get a non-negligible number of tau tridents, which is an important background to new physics searches. We identify promising kinematic features that may allow distinction of tau tridents from the usual neutrino charged-current background at DUNE, and thus could establish the observation of tau tridents for the first time. We also comment on the detection prospects at other accelerator and collider neutrino experiments.
@@ I Introduction
Of all SM particle species, the interactions of the tau neutrino () remain the least well-studied. The low cross sections for interactions, high tau-production thresholds and experimental difficulty to distinguish it from other neutrino states make studies of extraordinarily difficult. To date, the global dataset of directly observed interactions consists of only events at DONuT [3] and OPERA [4], making a future observation highly significant and a key target for the accelerator-neutrino programs.
@@
Even more challenging than simply observing charged-current (CC) interaction is to study neutrino trident production involving taus. Neutrino tridents are SM processes involving the creation of a charged-lepton pair via an energetic neutrino scattering with a nucleus, [5] (see Fig. 1). Although rare, these processes provide a precision test of the SM weak sector [6], as they are affected by both and -mediated interactions. They are also relevant for BSM physics searches, and in particular, tridents with will be an important background for charged lepton-flavor violation searches (see e.g., Refs. [7, 8, 9, 10, 11, 12]). So far, only the dimuon trident has been observed at level by CHARM [13] and CCFR [14] experiments (see also NuTeV [15, 16]) while the other final states have never been observed. This has motivated new proposals for observing tridents using future accelerator [17, 18, 19] and collider [20, 21] neutrino experiments, as well as at neutrino telescopes [22, 23]. Typically in the literature, studies have focused on the electron and muon tridents due to their relatively smaller energy threshold. The few existing studies [24, 17, 25] which briefly mentioned tau tridents have highlighted their low likelihood in standard accelerator-neutrino setups originating from scattering, particularly at near detectors.
@@
One key challenge for both CC interactions and tau tridents is the reconstruction of the tau lepton. One way to achieve this is by having excellent spatial resolution, at the micrometer level or below, which is typically attained using emulsion detectors such as FASER [26]. This allows to observe the vertex displacement from the tau propagation. Alternatively, detailed event reconstruction capabilities, as is the case for Liquid Argon Time Projection Chambers (LArTPCs) such as in DUNE [1], may allow for excellent signal-to-background ratios even if the displaced vertex from tau decay is not observed [27, 28].
@@
In this paper, we make the first detailed study of neutrino tridents involving one or two tau leptons at neutrino accelerator facilities, focusing specifically on DUNE Near Detector (ND) [29], at which the tau tridents were neglected in previous studies. We consider the DUNE standard CP-optimized flux, as well as the tau-optimized flux [30], for our calculation. We find non-negligible number of tau tridents at DUNE ND in both cases, and especially so in the tau-optimized mode. The prospective detection of these events is promising due to low intrinsic tau neutrino background and significantly different kinematic distributions between the tau trident signal and CC backgrounds. In the Appendix, we comment on trident events at current accelerator neutrino facilities and collider neutrino experiments.
II Neutrino trident cross section
@@
where for incoherent scattering is the nucleon mass GeV, and for coherent scattering is the nuclear mass which may lead to an enhancement of the latter scattering regime. The threshold for various trident channels are listed in Table 1 for three different nuclei, namely, Argon, Tungsten and Iron, which are respectively the detector materials for DUNE and FASER, as well as T2K and MINOS near detectors. The larger tau mass increases the energy threshold and reduces the likelihood of tau-production via the trident, but it remains possible if the neutrino energy reaches above threshold. Moreover, incoherent scattering for tau tridents has a much higher threshold than coherent scattering, unlike the electron and muon tridents, where both processes have similar thresholds since .
@@
Both coherent and incoherent scattering processes considered here involve elastic scattering of the neutrino off the target. The former is described by a nuclear form-factor, and the latter by a nucleon form factor. The cross sections may be calculated using [18]
@@
where denote the coherent and incoherent scattering regimes, respectively. is the photon virtuality, and , where is the four-momentum of the incoming neutrino. Eq. (II) factorizes the contributions to the trident process into transverse () and longitudinal () hadronic () and leptonic () components. The functions are dimensionless flux factors including the nuclear/nucleon form-factors , discussed in Appendix A. For the incoherent contribution to the cross section, Eq. (II) is further multiplied by a Pauli-blocking factor, derived by modeling the nucleus as an ideal Fermi Gas, assuming equal density of protons and neutrons [31],
@@
where is the spatial component of the momentum transfer to the nucleus, and is the Fermi momentum of the gas, taken to be MeV. The leptonic components are calculated as cross sections between a neutrino and an off-shell photon,
@@
where is the Lorentz invariant three-body phase space, and is the spin-averaged matrix element for , with a longitudinal or transversely polarized photon.
Here we remark on an alternative approach for calculating this component of the cross section, utilizing the equivalent photon approximation (EPA) [32]. To date, the only tau trident estimates for DUNE are given in Ref. [17], which employed the EPA, and deemed these processes inaccessible at DUNE energies. In the EPA, the infrared pole of the propagator for on-shell photon exchange between the lepton and nucleus is assumed to yield a dominant contribution to the cross section. As , the transverse cross section approaches that of an on-shell photon, and the longitudinal component vanishes. Thus, when applying the EPA the longitudinal contributions in Eq. (II) are omitted, and the total cross section scales roughly as [32, 17], where is the Fermi constant.
This EPA method overestimates the cross sections of the dielectron trident but is a good approximation for the coherent contribution to dimuon tridents [18, 25]. Ref. [18] establishes that EPA is generally not valid for tridents in regions where the virtual photon contributions are considerable. More relevant to our study, we note that the EPA will underestimate the cross section for tau tridents due to the larger contribution from off-shell photons.
@@
Our cross section results are shown in Fig. 2 for the three benchmark nuclei. We calculate the full scattering calculation using the formalism described in Refs. [18, 19]. Here we include both coherent (solid) and incoherent (dashed) neutrino scattering off nuclei. Following the methodology of Ref. [19], we estimate the cross section uncertainties for coherent (incoherent) scattering as follows. We assume a () uncertainty for the nuclear (nucleon) form factors. Higher order QED corrections scale as ; we assume a conservative () uncertainty for Argon, () uncertainty for Tungsten and () uncertainty for Iron. Our choice of weak mixing angle, , also leads to a () uncertainty. Additionally, given our simplified nuclear model of a Fermi gas, we include a conservative uncertainty for incoherent scattering. We add all our uncertainties in quadrature to get a total estimated uncertainty of () for Argon, () for Tungsten, and () for Iron. For further details, see Ref. [19].
@@
Due to lower neutrino energy threshold (see Table 1), the coherent contribution dominates the trident cross section at low neutrino energies. This is clearly evident for the tau trident processes for which there is a larger separation in the energy thresholds between the coherent and incoherent processes. As the neutrino energy increases, the minimum photon momentum-transfer, , necessary to generate the final state will decrease. As increases, the nuclear form factor decreases and the hadronic form factors come to dominate as the neutrino can resolve more of the nuclear structure (see Fig. 5 and 6 in Appendix A). This leads to an initial crossing point where the cross section is dominated first by the coherent scattering and then by incoherent. For larger incoming neutrino energies, the contribution from coherent scattering is proportional to (where is the atomic number), and the contribution from incoherent scattering scales with the number nucleons, i.e. proportional to . As the neutrino energy increases further, the enhancement from dependence of the coherent scattering will come to again dominate the evolution of the incoherent piece, leading to a second crossing. This is most prominent for the tau trident processes in Fig. 2.
@@
These two crossings are more pronounced in the tau trident case because they occur at much larger and more distinct neutrino energies. This is because of the presence of the tau, whose mass influences the momentum transfer . This pushes the crossing points to occur at larger neutrino energies and, therefore, there is a considerable region of energies sampled by the DUNE flux where incoherent production dominates the tau trident. We see from Fig 2 that coherent scattering dominates the muon trident much more rapidly as evolved with .
@@ III Neutrino tridents at DUNE
The relevant neutrino fluxes at various detectors considered here can be found in Appendix B. For DUNE, we consider the standard neutrino flux, as well as the tau-optimized configuration with more energetic neutrinos [39, 30].
In Fig. 3, we show the number of trident events, for all -initiated tau tridents as well as and , at DUNE ND with 67 ton of Argon and POT in the forward horn current, or neutrino mode (see Appendix C for antineutrino mode). The uncertainties shown come from the cross section uncertainties outlined in Sec. II. For simplicity, we do not take into account uncertainties in the neutrino fluxes, which can in principle be reduced with other measurements (e.g., neutrino-electron scattering) at DUNE ND. Since experimentally one may be able to distinguish between coherent and incoherent scattering by observing the hadronic activity, we show their corresponding number of events separately. For the ease of cross check with previous literature [18, 19], we report the dimuon trident events as well, although our main focus is on the taus. For instance, our event numbers for the dimuon tridents roughly match with those given in Ref. [19], after taking into account the differences in the flux and detector mass used there from the older DUNE configuration. As for the tau tridents, in total, we expect () events in the standard mode and () events in the tau-optimized mode for the () trident. Most of these events come from incoherent scattering due to DUNE’s energy spectrum. This is in sharp contrast with the dimuon trident, where the coherent contribution is the dominant one. The tau-optimized mode gives more events because the flux peaks at higher energies, making it easier to pass the energy threshold for the tau tridents. For ditau tridents, we expect less than one event even in the tau-optimized mode. For completeness, we also report the numbers for and as well as their antineutrino counterparts in Fig. 3.
In Appendix C, we also show the number of trident events for FASER, T2K INGRID near detector, MINOS/MINOS+.
@@ IV Tau Identification from Tridents at DUNE
Detecting tau leptons in a realistic neutrino experimental setup like DUNE is notoriously challenging. First of all, LArTPCs cannot identify taus on an event-by-event basis due to the very short displaced vertex. Besides, the tau production cross section, either from intrinsic beam or tridents, is very low, making virtually any experimental setup statistics-limited for tau observation. Tau reconstruction suffers from large backgrounds induce by the copious beam component. Misreconstruction of final state particles could significantly degrade the signal-to-background ratio. This has been considered a major challenge in detection at DUNE in the context of BSM studies [35, 36, 40, 37]. However, recent developments based on advanced machine-learning algorithms hold some promise [41, 42, 43, 44].
@@
The basic idea is to separate the tau signal events from backgrounds using kinematic differences, similar to those used in NOMAD [45]. Reference [44] ranked six kinematic variables providing the highest signal-to-background for tau reconstruction. The ratio of the transverse momenta, , was identified as the rank-1 discriminator. Note that is inferred from the total visible transverse momentum, including the muon, charged and neutral pions from tau decays, and possibly protons in the case of incoherent scattering. This ratio is particularly relevant for our tau trident signal which always comes with additional missing energy from neutrinos, and therefore, we expect a larger for the signal compared to the background.
In Fig. 4, we show the distributions for the trident with incoherent (left) and coherent (right) contributions in blue for two benchmark average neutrino energies derived from their event distributions (see Fig. 12 in Appendix D). Here we consider the hadronic decays of tau, due to their larger branching ratio (65%), compared to the leptonic decays (17%–18%). We have simulated the hadronic tau decays using PYTHIA 8.3 [46]. For comparison, we show the dominant background from CC extracted from the works of Aurisano-Rajaolisoa-Razafinime-Sousa (ARRS) [41, 42, 43, 44], which used the GENIE event generator [47]. We also have generated the same background using the NuWro event generator [48]. Although the GENIE and NuWro distributions differ by 20%, they both shrink at larger where most of the signal is. While not a full analysis, this serves to illustrate the point that discriminators like could mitigate the large CC background, enhance the signal-to-background ratio, and thus potentially allow us to observe the handful of tau trident events at DUNE ND. A detailed experimental feasibility study is left as future work.
For completeness, we also present the tau trident event rates for MINOS/MINOS+, T2K INGRID and FASER/FASER2 in Appendix C.

### v2

@@ Abstract
We present the first detailed study of Standard Model neutrino tridents involving tau leptons at the near detectors of accelerator neutrino facilities. The rates of these processes were previously thought to be negligible, even at future facilities. Our full calculation, including both coherent and incoherent scatterings, reveals that the DUNE near detector could observe a considerable number of tau tridents – an important background to new physics searches. We identify promising kinematic features that may allow distinction of tau tridents from the usual neutrino charged-current background at DUNE, and thus establish the observation of tau tridents for the first time. We also comment on the detection prospects at other accelerator and collider neutrino experiments.
@@ I Introduction
Of all SM particle species, the tau neutrino () remains the least well-studied. The low cross sections for interactions, high tau-production thresholds, and experimental difficulty of distinguishing it from other neutrino states make studies of extraordinarily difficult. To date, the global dataset of directly observed interactions consists of only events at the DONuT [3] and OPERA [4] experiments, making future observation highly significant and a key target for accelerator-neutrino programs.
@@
Even more challenging than simply observing charged-current (CC) interactions is to study neutrino trident production involving taus. Neutrino tridents are SM processes involving the creation of a charged-lepton pair via an energetic neutrino scattering with a nucleus, [5] (see Fig. 1). Although rare, these processes provide a precision test of the SM weak sector, as they involve both and -mediated interactions (see e.g., [6, 7, 8, 9, 10]). They are also relevant for BSM physics searches, and in particular, tridents with will be an important background for charged lepton-flavor violation searches (see e.g., [11, 12, 13, 14, 15, 16]). So far, only the dimuon trident has been observed at level by CHARM [7] and CCFR [8] experiments (see also NuTeV [17, 9]), whereas the other final states are yet to be observed. This has motivated new proposals for observing tridents using future accelerator [18, 19, 20] and collider [21, 22] neutrino experiments, as well as at neutrino telescopes [23, 24]. Typically in the literature, studies have focused on the electron and muon tridents due to their relatively smaller energy threshold. The few existing studies [25, 18, 26] which briefly mentioned tau tridents have highlighted their low likelihood in standard accelerator-neutrino setups originating from scattering, particularly at near detectors.
@@
A key challenge for both CC interactions and tau tridents is the reconstruction of the tau lepton. One way to achieve this is by having excellent spatial resolution, at the micrometer level or below, which is typically attained using emulsion detectors, such as those at FASER [27]. This allows observation of the vertex displacement from tau propagation. Alternatively, detailed event reconstruction capabilities, as is the case for Liquid Argon Time Projection Chambers (LArTPCs) such as in DUNE [1], can lead to excellent signal-to-background ratios even if the displaced vertex from tau decay cannot be resolved [28, 29].
In this work, we perform the first detailed study of neutrino tridents involving one or two tau leptons at
neutrino accelerator facilities, focusing specifically on the DUNE Near Detector (ND) [30], at which the tau tridents were neglected in previous studies.
@@
We consider the DUNE standard CP-optimized flux, as well as the tau-optimized flux [31], for our calculation. We find non-negligible number of tau tridents at DUNE-ND in both cases, and especially so in the proposed tau-optimized mode [32]. We also highlight the different behavior of the tau-containing trident cross sections due to a non-negligible final-state charged-lepton mass, in contrast with trident processes with lighter leptons better studied for accelerator facilities in the literature. The prospective detection of events is promising due to low intrinsic tau neutrino background and significantly different kinematic distributions between the tau trident signal and CC backgrounds.
The rest of the paper is organized as follows: In Section II, we calculate the trident production cross sections. In Section III, we estimate the number of trident events at DUNE. In Section IV, we discuss how tau tridents can be identified at DUNE. Our conclusions are given in Section V. In Appendix A, we remark on the Equivalent Photon Approximation (EPA) approach to trident calculations. In Appendix B, we outline the nuclear and nucleon form factors. In Appendix C, we show the neutrino fluxes used in our calculation. In Appendix D, we give the number of trident events at other accelerator/collider neutrino experiments. Finally, Appendix E shows the tau trident event distributions at DUNE.
II Neutrino trident production
@@
where for incoherent scattering is the nucleon mass GeV, and for coherent scattering is the nuclear mass which may lead to an enhancement of the latter scattering regime. The threshold for various trident channels are listed in Table 1 for three different nuclei, namely, argon, tungsten and iron, which are respectively the detector materials for DUNE, FASER, and the T2K and MINOS near detectors. The larger tau mass increases the energy threshold and reduces the likelihood of tau-production via the trident, but it remains possible if the neutrino energy reaches above threshold. Moreover, incoherent scattering for tau tridents has a much higher threshold than coherent scattering, unlike the electron and muon tridents, where both processes have similar thresholds since .
@@
II.1 Cross section calculation
We calculate the full elastic scattering for both coherent and incoherent processes using the procedure described in Ref. [19], where
@@ II.1 Cross section calculation
where denote the coherent and incoherent scattering regimes, respectively. is the photon virtuality, and , where is the four-momentum of the incoming neutrino. Eq. (2) factorizes the contributions to the trident process into transverse () and longitudinal () hadronic () and leptonic () components. The functions are dimensionless flux factors including the nuclear (nucleon) form-factors for coherent (incoherent) scattering, as will be discussed below.
@@
The leptonic components are calculated as cross sections between a neutrino and an off-shell photon,
@@
where is the Lorentz invariant three-body phase space, and is the spin-averaged matrix element for , with a longitudinal or transversely polarized photon. We comment on the alternative Equivalent Photon Approximation (EPA) method for calculating in Appendix A.
II.1.1 Coherent scattering
II.1.2 Incoherent scattering
In the incoherent (diffractive) regime, the flux factors appearing in Eq. (2) have the form
where is the proton or neutron mass, and are the nucleon form factors, defined as
where . The functions and appearing in the nucleon flux factors via Eq. (10) and (11) can be related to the electric and magnetic form-factors of the nucleons as
where the form-factors used in this work are outlined explicitly in Appendix B.
For this incoherent contribution to the cross section, Eq. (2) is further multiplied by a Pauli-blocking factor, derived by modeling the nucleus as an ideal Fermi Gas, assuming equal density of protons and neutrons,
where is the spatial component of the momentum transfer to the nucleus, and is the Fermi momentum of the gas, taken to be MeV [33].
II.2 Total cross sections for benchmark nuclei
Our cross section results are shown in Fig. 2 for a number of benchmark nuclei: where argon is relevant for DUNE, tungsten and iron for other accelerator neutrino facilities: FASER, T2K INGRID near detector and MINOS/MINOS+ for contrast; see Appendix D.
@@ II.2 Total cross sections for benchmark nuclei
Following the methodology of Ref. [20], we estimate the cross section uncertainties for coherent (incoherent) scattering as follows. We assume a () uncertainty for the nuclear (nucleon) form factors.111This is the error budget for a normalization uncertainty. If we consider an uncertainty in the shape of the form factor, we do not expect significant impact on our final results as the missing-energy nature of the neutrino final state will smear any potential effects. Higher order QED corrections scale as ; we assume a conservative () uncertainty for argon, () uncertainty for tungsten and () uncertainty for iron. Our choice of weak mixing angle, , also leads to a () uncertainty. Additionally, given our simplified nuclear model of a Fermi gas, we include a conservative uncertainty for incoherent scattering. We add all our uncertainties in quadrature to get a total estimated uncertainty of () for argon, () for tungsten, and () for iron.
@@
Due to lower neutrino energy threshold (see Table 1), the coherent contribution dominates the trident cross section at low neutrino energies. This is clearly evident for the tau trident processes for which there is a larger separation in the energy thresholds between the coherent and incoherent processes. As the neutrino energy increases, the minimum photon momentum-transfer, , necessary to generate the final state will decrease. As increases, the nuclear form factor decreases and the hadronic form factors come to dominate as the neutrino can resolve more of the nuclear structure (see Appendix B). This leads to an initial crossing point where the cross section is dominated first by the coherent scattering and then by incoherent. For larger incoming neutrino energies, the contribution from coherent scattering is proportional to (where is the atomic number), and the contribution from incoherent scattering scales with the number nucleons, i.e. proportional to . As the neutrino energy increases further, the enhancement from the dependence of the coherent scattering will again dominate over the evolution of the incoherent piece, leading to a second crossing. This is most prominent for the tau trident processes in Fig. 2.
@@
These two crossings are more pronounced in the tau trident case because they occur at much larger and more distinct neutrino energies. This is because of the presence of the tau, whose mass influences the momentum transfer . This pushes the crossing points to occur at larger neutrino energies and, therefore, there is a considerable region of energies sampled by the DUNE flux where incoherent production dominates the tau trident. We see from Fig. 2 that coherent scattering dominates the muon trident much more rapidly as evolved with .
@@ III Neutrino tridents at DUNE
Here is the trident cross section. The relevant neutrino fluxes at various detectors considered here can be found in Appendix C. For DUNE, we consider the standard neutrino flux, as well as the tau-optimized configuration with more energetic neutrinos [32, 31].
In Fig. 3, we show the number of trident events, for all -initiated tau tridents as well as and , at DUNE-ND with 67 ton of argon and POT in the forward horn current, or neutrino mode (see Appendix D for antineutrino mode). The uncertainties shown come from the cross section uncertainties outlined above. For simplicity, we do not take into account uncertainties in the neutrino fluxes, which can in principle be reduced with other measurements (e.g., neutrino-electron scattering) at DUNE-ND. Since experimentally one may be able to distinguish between coherent and incoherent scattering by observing the hadronic activity, we show their corresponding number of events separately. For the ease of cross check with previous literature [19, 20], we report the dimuon trident events as well, although our main focus is on the taus. For instance, our event numbers for the dimuon tridents roughly match with those given in Ref. [20], after taking into account the differences in the flux and detector mass used there from the older DUNE configuration. As for the tau tridents, in total, we expect () events in the standard mode and () events in the tau-optimized mode for the () trident running in the forward-horn configuration. Most of these events come from incoherent scattering due to DUNE’s energy spectrum. This is in sharp contrast with the dimuon trident, where the coherent contribution is the dominant one. The tau-optimized mode gives more events because the flux peaks at higher energies, making it easier to pass the energy threshold for the tau tridents. For ditau tridents, we expect less than one event even in the tau-optimized mode. For completeness, we also report the numbers for and as well as their antineutrino counterparts in Fig. 3. In Appendix D we also show the number of trident events for FASER, T2K INGRID near detector and MINOS/MINOS+ for comparison.
@@ IV Tau Identification from Tridents at DUNE
Detecting tau leptons in a realistic neutrino experimental setup like DUNE is notoriously challenging. First of all, LArTPCs cannot identify taus on an event-by-event basis due to the very short displaced vertex. Besides, the tau production cross section, either from intrinsic beam or tridents, is very low, making virtually any experimental setup statistics-limited for tau observation. Tau reconstruction suffers from large backgrounds induced by the copious beam component. Misreconstruction of final state particles could significantly degrade the signal-to-background ratio. This has been considered a major challenge in detection at DUNE in the context of BSM studies [36, 37, 40, 38]. However, recent developments based on advanced machine-learning algorithms hold some promise [41, 42, 43, 44].
@@
The basic idea is to separate the tau signal events from backgrounds using kinematic differences, similar to those used in NOMAD [45]. Reference [44] ranked six kinematic variables providing the highest signal-to-background for tau reconstruction. The ratio of the transverse momenta, , was identified as the rank-1 discriminator. This ratio is particularly relevant for our tau trident signal which always comes with additional missing energy from neutrinos, and therefore, we expect a larger for the signal compared to the background.
In Fig. 4, we show the distributions for the trident with incoherent (left) and coherent (right) contributions in blue for two benchmark average neutrino energies derived from their event distributions (Fig. 12 in Appendix E). Here we consider the hadronic decays of tau, due to their larger branching ratio (65%), compared to the leptonic decays (35%). We have simulated the hadronic tau decays using PYTHIA 8.3 [46] and subsequently boosted them to our trident frame. For comparison, we show the dominant background from CC extracted from the works of Refs. [41, 42, 43, 44], which used the official GENIE/GEANT4-based DUNE simulation [47, 48, 49]. We also have generated the same background using the NuWro event generator [50]. To account for detector effects, we smeared the 3-momenta and polar angles of particles as follows. For coherent scattering off argon, we smear by on each component. For incoherent scattering, is inferred instead from the visible outgoing particles (we treat neutrons as invisible). We assume a 2% (10%) energy and () angular smearing for muons (other particles).
Although the DUNE and NuWro distributions differ by 20%, they both shrink at larger where most of the signal is. To estimate the impact of as a discriminator, we generate distributions for neutrino energies ranging from . We then applied an cut and extracted the percentage of signal events that pass this cut at a given , producing an energy-dependent efficiency. Including this efficiency, the expected number of coherent scattering trident events at DUNE ND in neutrino mode changed from to (standard beam) and from to (tau-optimized beam). For incoherent scattering, the corresponding numbers changed from to and from to . While not a full analysis, this serves to illustrate the point that discriminators like could mitigate the large CC background, enhance the signal-to-background ratio, and thus potentially allow us to observe the handful of tau trident events at DUNE-ND. A detailed experimental feasibility study is left as future work. Moreover, high-energy and high-intensity neutrino beams available at future colliders could provide another promising avenue for tau tridents; see e.g. Ref. [51].